OpenStudy (erikaxx):

another linear differential equation question, pls help will give medal & fan forevs

OpenStudy (erikaxx):

@phi

OpenStudy (phi):

do you really mean the exponent is 2-6 ?

OpenStudy (erikaxx):

yes that the exponent written in the book

OpenStudy (phi):

that's the same as r^-4 ??

OpenStudy (erikaxx):

yes

OpenStudy (erikaxx):

so what to do now

OpenStudy (phi):

it looks like a typo and they meant \[ (4r ^{2}-6) dr +r^3 ds = 0 \]

OpenStudy (erikaxx):

oops the 6 should be in exponent

OpenStudy (erikaxx):

\[(4r ^{-4}) dr + r^3 ds = 0\]

OpenStudy (erikaxx):

whats next

OpenStudy (phi):

I would multiply the equation by r^-3

OpenStudy (erikaxx):

oh so it will be \[4r ^{-7} dr + \frac{ 1 }{ ds } = 0\]

OpenStudy (erikaxx):

then?

OpenStudy (phi):

how did you flip ds ?

OpenStudy (erikaxx):

oh just DS only

OpenStudy (erikaxx):

@phi then what now pls

OpenStudy (phi):

you can move ds to the other side \[ 4 r^{-7} \ dr = - ds \] now integrate both sides

OpenStudy (erikaxx):

\[\frac{ 4r ^{-8} }{ 8 } = -ds\]

OpenStudy (erikaxx):

\[\frac{ r ^{-8} }{ 2 } = -ds\] @phi

OpenStudy (phi):

the right side would become - s + C you have to do the same thing (i.e. integrate) to both sides to keep it an equation

OpenStudy (phi):

also, the left side should be negative, right ?

OpenStudy (erikaxx):

\[\frac{ r ^{-8} }{ 2 } - s = c\]

OpenStudy (phi):

so I would have done \[ -\frac{ r ^{-8} }{ 2 } = - s + c\] or multiplying by -1: \[ \frac{ r ^{-8} }{ 2 }= s+ c\] (-c is still just a constant, all it c) and finally \[ \frac{ r ^{-8} }{ 2 } - s = c\]

OpenStudy (erikaxx):

@phi but the final answer in book is \[s = \frac{ 3 }{ r^2 }+ \frac{ c }{ r^4 }\]

OpenStudy (phi):

which makes me think the original question has a typo in it.

OpenStudy (erikaxx):

ohh

OpenStudy (loser66):

Can you take a snapshot of the problem?

OpenStudy (erikaxx):

okay

OpenStudy (phi):

it looks like the original problem should be \[ (4r^{-2} +6 )\ dr + r^3 \ ds =0\]

OpenStudy (erikaxx):

oh, lets try it

OpenStudy (erikaxx):

so we divide it by?

OpenStudy (phi):

yes, divide by r^3, and move ds to the other side

OpenStudy (erikaxx):

\[(4r + \frac{ 6 }{ r^3 }) dr + ds = 0\]

OpenStudy (phi):

horribly wrong

OpenStudy (erikaxx):

4/r

OpenStudy (erikaxx):